# Diffusion approximation of the stationary distribution of a two-level single server queue

### 18 Dec 2023 v1

@@ 3.2 Main results
=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{2}\not=0.\end{cases}$ (3.13)
@@ 5 Concluding remarks
In this section, we discuss two subjects, a process limit for the 2-level $GI/G/1$ queue in heavy traffic and a verification of (3.11) in Remark 3.1.
@@ 5.1 Process limit
Recall that one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation:
@@
which follows from (4.3), (4.54) and (4.63) since $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$.
We like to identify this stationary distribution $\widetilde{\nu}$ by that of the reflecting diffusion process. Assume that the diffusion scaled process $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to a nonnegative valued stochastic process $Z(\cdot)\equiv\{Z(t);t\geq 0\}$ as $r\downarrow 0$, and define real-valued functions $\sigma(x)$ and $b(x)$ for $x\in\mathbb{R}_{+}$ as
@@
However, $Y(\cdot)$ may not be absolutely continuous. So far, assume that there is a weak solution of (5.2). For notational convenience, we continue to use the same notations for this weak solution. To apply Ito integration rule to (5.2), let us compute the quadratic variation $\big{[}Z\big{]}_{t}$ of $Z(\cdot)$. Define $M(\cdot)\equiv\{M(t);t\geq 0\}$ as
@@
then $M(\cdot)$ is a martingale. Denote the quadratic variations of $Z(\cdot)$ and $M(\cdot)$, respectively, by $\big{[}Z\big{]}_{t}$ and $\big{[}M\big{]}_{t}$. Then, it follows from (5.2) that
@@
Hence, by Ito integration rule, we have, for $f\in C^{2}_{b}(\mathbb{R}_{+})$, where $C^{2}_{b}(\mathbb{R}_{+})$ is the set of all bounded and twice differentiable function from $\mathbb{R}_{+}$ to $\mathbb{R}$ which have the derivatives from the right at the origin, we have
@@
where $\Gamma$ is operator on the space $C^{2}_{b}(\mathbb{R}_{+})$ defined as
$\Gamma f(x)=-b(x)f^{\prime}(x)+\frac{1}{2}\sigma^{2}(x)f^{\prime\prime}(x),\qquad x\geq 0.$
In particular, let $f$ be
$f(x)=e^{\theta_{1}x}1(0\leq x\leq\ell_{0})+e^{\theta_{2}x}1(x>\ell_{0}),\qquad\theta_{i}\leq 0,i=1,2.$ (5.6)
If $Z(\cdot)$ is a stationary process, then taking the expectation of (5.5) for this $f$ yields
$\Big{(}\!-b_{1}\theta_{1}+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta_{1}^{2}\Big{)}\widetilde{\varphi}_{1}(\theta_{1})+\Big{(}\!-b_{2}\theta_{2}+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta_{2}^{2}\Big{)}\widetilde{\varphi}_{2}(\theta_{2})+\theta_{1}\widetilde{\varphi}_{1}(0)\mathbb{E}[Y(1)]=0,$
which implies that $\widetilde{\varphi}_{1}(\theta_{1})$ is the moment generating function of the exponential distribution and $\widetilde{\varphi}_{2}(\theta_{2})=0$ for all $\theta_{2}\leq 0$. Obviously, this is absurd. Thus, there is no solution $Z(\cdot)$ for (5.2) if $Z(\cdot)$ is a stationary process.
What is a right stochastic integral equation for $Z(\cdot)$ Although $Z(\cdot)$ is not reflected at $\ell_{0}$, we may consider the discontinuous change of $\sigma(Z(u))$ at $Z(u)=\ell_{0}$ as something like reflection. This motivates us to define local times $L_{\ell_{0}}^{-}(t)$ and $L_{\ell_{0}}^{+}(t)$ at $\ell_{0}$ (e.g. see Theorem 22.5 of [6] for the definition of a local time) by
@@
and replace (5.2) by
@@
Then, applying Ito integration rule to this stochastic integral equation for $f$ of (5.6) and taking for $Z(\cdot)$ to be stationary yield (5.1). This suggests the following conjectures.
@@
    $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to $Z(\cdot)$ satisfying (5.9) as $r\downarrow 0$.
@@
    Stochastic integral equation (5.9) has a unique weak solution.
Both are hard to verify, so we leave them for future work. Since $Z(\cdot)$ of (5.9) is specified by $b_{i},c_{i},\sigma_{i}^{2}$ for $i=1,2$ and $\ell_{0}$, its stationary distribution must be $\widetilde{\nu}$ of (3.7). This supports the conjecture in Remark 3.1 because of (ii) of Theorem 3.1.

### 10 Feb 2024 v3

@@ Abstract
Note: This version corrects Section 5.1 in addition to fixing typos.
@@ 3.2 Main results
=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.13)
@@ 5 Concluding remarks
In this section, we intuitively answer two questions, what is a process limit for the 2-level $GI/G/1$ queue in heavy traffic and how to verify (3.11) in Remark 3.1.
@@ 5.1 Process limit
Recall that one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation, that is, BAR:
@@
which follows from (4.3), (4.54) and (4.63) since $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$. For $b_{1}=0$, BAR is obtained from (5.1) by letting $b_{1}\downarrow 0$. So, we only consider the case that $b_{1}=0$.
We like to identify this stationary distribution $\widetilde{\nu}$ by that of a reflecting diffusion process. Assume that the diffusion scaled process $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to a nonnegative valued stochastic process $Z(\cdot)\equiv\{Z(t);t\geq 0\}$ as $r\downarrow 0$. We aim to characterize this $Z(\cdot)$ by a stochastic integral equation. To this end, define real-valued functions $\sigma(x)$ and $b(x)$ for $x\in\mathbb{R}_{+}$ as
@@
However, $Y(\cdot)$ is nondecreasing, but may not be absolutely continuous. So far, we just assume that there is a weak solution of (5.2). For notational convenience, we continue to use the same notations for this weak solution. To apply Ito integration rule to (5.2), let us compute the quadratic variation $\big{[}Z\big{]}_{t}$ of $Z(\cdot)$. Define $M(\cdot)\equiv\{M(t);t\geq 0\}$ as
@@
then $M(\cdot)$ is a martingale. Denote the quadratic variations of $Z(\cdot)$ and $M(\cdot)$, respectively, by $\big{[}Z\big{]}_{t}$ and $\big{[}M\big{]}_{t}$. Then, $Z(t)$ and $Y(t)$ are continuous in $t$, so it follows from (5.2) that
@@
We next define local time $L_{x}(t)$ for $x\geq 0$ as
@@
where ${\rm sgn}(x-)=1(x>0)-1(x\leq 0)$ (e.g., see (4) of Chapter 22 of [6]). Note that, for any measurable function $f\geq 0$,
[…]
by Theorem 22.5 of [6]. Note that $L_{x}(t)$ is $a.s.$ continuous in $t\geq 0$, but may not be so in $x$ if $Z(\cdot)$ is not a martingale. So, let $L^{+}_{x}(t)$ and $L^{-}_{x}(t)$ be the right and left continuous versions on $x$, respectively, of $L_{x}(t)$ for each fixed $t>0$. Namely, $L^{+}_{x}(t)=\lim_{\varepsilon\downarrow 0}L_{x+\varepsilon}(t)$ and $L^{-}_{x}(t)=\lim_{\varepsilon\downarrow 0}L_{x-\varepsilon}(t)$. As we will see, their difference is a key for our analysis.
To derive a BAR for (5.2), we use the extended Ito formula for a convex test function $f$ for the local time $L^{-}_{x}(t)$, which is given as
[…]
where $f^{\prime}_{+}(x)=f^{\prime}(x+)$, that is, left derivative of $f$ at $x$, and $\mu_{f^{\prime}_{+}}$ is a measure on $(\mathbb{R}^{+},{\mathcal{B}}(\mathbb{R}_{+}))$ defined by
$\mu_{f^{\prime}_{+}}((x,y])=f^{\prime}_{+}(x)-f^{\prime}_{+}(y),\qquad x<y\mbox{ in }\mathbb{R}_{+}.$
Note that (5.7) is the left-continuous local time counterpart of Theorem 22.5 of [6], which takes the right-continuous local time $L^{+}_{x}(t)$ and $\mu_{f^{\prime}_{-}}$ instead of $L^{-}_{x}(t)$ and $\mu_{f^{\prime}_{+}}$.
For (5.7), we take the test function $f$ with parameter $\theta\geq 0$, which is defined as $f(x)=e^{\theta h^{-}_{\ell_{0}}(x)}$, where $h^{-}_{\ell_{0}}(x)=(\ell_{0}-x)1(0\leq x\leq\ell_{0})$. Then, $f$ is a convex function for $x\geq 0$, and $f^{(n)}_{+}(x)=\theta^{n}e^{\theta x}1(0\leq x<\ell_{0})$ for $n\geq 1$. Since $\mu_{f^{\prime}_{+}}(\{\ell_{0}\})=-f^{\prime}(\ell_{0}-)$ and $\mu_{f^{\prime}_{+}}$ has the density $f^{\prime\prime}_{+}(x)=\theta^{2}e^{\theta x}1(0\leq x<\ell_{0})$, it follows from (5.4) and (5.6) that
[…]
$=\frac{1}{2}\theta^{2}\int_{0}^{t}e^{\theta Z(u)}1(0\leq Z(u)<\ell_{0})d[Z]_{u}-\frac{1}{2}\theta e^{\theta\ell_{0}}L^{-}_{\ell_{0}}(t).$
Hence, the extended Ito formula (5.7) becomes
@@
If $Z(\cdot)$ is a stationary process, then taking the expectation of (5.1) for $t=1$ yields
$\Big{(}-b_{1}\theta+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta^{2}\Big{)}\widetilde{\varphi}_{1}(\theta)-\frac{1}{2}\theta e^{\theta\ell_{0}}\mathbb{E}[L^{-}_{\ell_{0}}(1)]+\theta\widetilde{\varphi}_{1}(0)\mathbb{E}[Y(1)]=0,$ (5.9)
where $\widetilde{\varphi}_{1}(\theta)$ is the moment generating function of $Z(1)$ on $0\leq Z(1)\leq\ell_{0}$.
Similarly, let $\widetilde{\varphi}_{2}(\theta)$ be the moment generating function of $Z(1)$ on $Z(1)>\ell_{0}$, then, taking the right-continuous version $L^{+}_{x}(t)$ of the local time $L_{x}(t)$, and letting $f(x)=-e^{\theta h^{+}_{\ell_{0}}}$ for $\theta\leq 0$, where $h^{+}_{\ell_{0}}=(x-\ell_{0})1(x>\ell_{0})$, we have
$\Big{(}-b_{2}\theta+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta^{2}\Big{)}\widetilde{\varphi}_{2}(\theta)+\frac{1}{2}\theta e^{\theta\ell_{0}}\mathbb{E}[L^{+}_{\ell_{0}}(1)]=0.$ (5.10)
We finally note that
@@
[…]
Then, combining (5.9) and (5.10), we arrive at (5.1). Thus, the BAR of $Z(\cdot)$ is identical with (5.1). Based on this fact, we conjecture that
@@
    there is a unique weak solution $Z(\cdot)$ for the stochastic integral equation (5.2),
@@
    $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to the $Z(\cdot)$ of (a),
  • (c)
    The $\widetilde{\nu}$ of (3.7) in Theorem 3.1 is the unique solution of the BAR (5.1) and the stationary distribution of $Z(t)$.
These conjectures are challenging, but we expect that they are verifiable.